News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Friday 6 August 2021 9.00am EDT

During the last week Chandra completed the observing schedule as planned.

Real-time procedures were executed Jul 30 - Aug 2 to dump OBC-A and -B memory as a follow-up to the patches uplinked on Jul 30. The dumps will be used to update the baseline memory images maintained on the ground. Another real-time procedure was executed on Aug 1 to swap the prime solid-state recorder (SSR) from the A-side to the B-side. This swap concludes a six-month interval of health and trending data collection on SSR-A. Additional real-time procedures were executed on Aug 4 to perform a routine self-check of the Electrical Interface Assembly (EIA) Sequencer and to perform a diagnostic dump of the Control Processing Electronics (CPE).

A Chandra image release was issued on Aug 5 to present data on the black hole in V404 Cygni from Chandra and from other telescopes. V404 Cygni is a binary system in which a less massive star orbits the black hole leading to occasional bursts of X-rays. This burst created high-energy rings, a phenomenon known as light echoes. The light echoes around V404 Cygni were produced when the X-rays from the black hole system bounced off of dust clouds between V404 Cygni and Earth. The rings can be used to understand information about not only about the black hole's behavior, but also about the material between V404 Cygni and Earth. For more details see:

The schedule of targets for the next week is shown below and includes an observation of IGR J16318-4848 coordinated with NuSTAR, and observations of GRB210726A, which are follow-up to a Target of Opportunity accepted on Jul 27.

        Radiation Belts                     Aug  8
        SPT-CLJ2031-4037       ACIS-S
        PSRJ0821-4300          HRC-S /LETG
        MR2251-178             HRC-S /LETG
        GRB210726A             ACIS-S       Aug 10
        2RXSJ201634.6-07114    ACIS-S
        GRB210726A             ACIS-S
        4U1916-053             ACIS-S/HETG  Aug 11
        Radiation Belts
        4U1916-053             ACIS-S/HETG
        PSRJ0821-4300          HRC-S /LETG
        MR2251-178             HRC-S /LETG  Aug 12
        IC342SW                ACIS-I
        hetej1900.1-2455       ACIS-S       Aug 13
        IGRJ16318-4848         ACIS-S/HETG
        Radiation Belts                     Aug 14
        Westerlund1            ACIS-I
        4U1916-053             ACIS-S/HETG
        Westerlund1            ACIS-I
        2RXSJ211402.0-29395    ACIS-S       Aug 15
        HD25874                HRC-I


All spacecraft subsystems continued to support nominal operations.

Return to Status Reports